OpenStudy (deepolisnoob):

I have 2 questions: 1.) Given f(x)=-6x^2, g(x)=-7x+2, and h(x)=(square root of x), find [(f+g) of h](x). 2.) Let f(x)=6x+2. Find (f(x)-f(a))/(x-a), if x cannot equal a. I believe the first question's answer is -6x-7(square root of x)+2, but I don't understand the second question

OpenStudy (aum):

The answer to the first question is correct: \(-6x -7\sqrt{x}+2\). 2) \[ f(x) = 6x + 2 \\ \frac{f(x)-f(a)}{x-a} = \frac{(6x + 2) - (6a+2)}{x-a} = \frac{6x + 2 - 6a-2}{x-a} = \frac{6(x - a)}{x-a} =6\\ \]

OpenStudy (deepolisnoob):

How did you plug in the value of f(a)? Do you just copy the f(x) value, replacing x with a?